### v1

# Mamba4Rec: Towards Efficient Sequential Recommendation with Selective State Space Models

Abstract.
Sequential recommendation aims to estimate the dynamic user preferences and sequential dependencies among historical user behaviors. Although Transformer-based models have proven to be effective for sequential recommendation, they suffer from the inference inefficiency problem stemming from the quadratic computational complexity of attention operators, especially for long-range behavior sequences. Inspired by the recent success of state space models (SSMs), we propose Mamba4Rec, which is the first work to explore the potential of selective SSMs for efficient sequential recommendation. Built upon the basic Mamba block which is a selective SSM with an efficient hardware-aware parallel algorithm, we incorporate a series of sequential modeling techniques to further promote the model performance and meanwhile maintain the inference efficiency. Experiments on two public datasets demonstrate that Mamba4Rec is able to well address the effectiveness-efficiency dilemma, and defeat both RNN- and attention-based baselines in terms of both effectiveness and efficiency. The code is available at https://github.com/chengkai-liu/Mamba4Rec.
Sequential Recommendation, State Space Models
1. Introduction
Personalized online services rely heavily on sequential recommender systems to capture dynamic user preferences (wang2021survey; lu2015recommender; lin2023map). These systems aim to predict users’ future interactions by effectively modeling sequential dependencies among historical user behaviors. Earlier neural approaches adopt Convolutional Neural Networks (CNNs) (tang2018personalized) and Recurrent Neural Networks (RNNs) (hidasi2015session; li2017neural) which have pioneered the application of neural networks in sequential recommendation yet suffer from the catastrophic forgetting issue (schak2019study; kirkpatrick2017overcoming). More recently, researchers start to introduce Transformer (vaswani2017attention) as the backbone module and leverage the self-attention mechanism to model the dynamic user behavior sequences (kang2018self; yuan2022multi). In spite of their remarkable performance, these attention-based methods generally suffer from the inference inefficiency problem due to the quadratic computational complexity inherent in attention operators, especially for long user behavior sequences or large-scale datasets.
To address the dilemma of recommendation performance and inference efficiency, in this paper, we consider the state space model (SSM) (gu2021efficiently) as the core operator for sequential recommendation. SSMs (e.g., S4 (gu2021efficiently), S5 (smith2022simplified), Mega (ma2022mega)) have been widely adopted as alternatives to RNNs, CNNs, and Transformers in various natural language tasks. They inherently possess inference efficiency due to their recurrent nature, and meanwhile enjoy strong performance to handle long-range dependencies through structured state matrices. Moreover, a recently proposed SSM variant (i.e., Mamba (gu2023mamba)) further introduces the data-dependent selective mechanism with efficient hardware-aware design to tackle the data- and time-invariant issue of prior SSMs. This allows the model to selectively extract essential knowledge and filter out irrelevant noise according to the input data, thus leading to superior sequential modeling performance.
In this work, we propose Mamba4Rec, which is the first work to explore the potential of selective SSM for efficient sequential recommendation. Built upon the basic Mamba block (gu2023mamba), we incorporate and discuss a series of techniques (e.g., position embedding (vaswani2017attention), residual connection (he2016deep), layer normalization (ba2016layer), and position-wise feed-forward network (vaswani2017attention)), which supplements the system non-linearity, stabilizes the training dynamic, and thereby further promotes the sequential modeling capability.
The main contributions of this paper are as follows:
  • We are the first to explore the potential of selective SSMs for sequential recommendation, which addresses the dilemma of recommendation performance and inference efficiency.
  • We propose Mamba4Rec by incorporating the basic Mamba block with various techniques like layer normalization and feed-forward networks, which further improves the sequential modeling capability without sacrificing the inference efficiency.
  • Experiments on two public datasets demonstrate the superiority of our proposed Mamba4Rec compared with RNN- and attention-based baselines in terms of both effectiveness and efficiency.
2. Preliminaries
2.1. Problem Statement
In sequential recommendation, let $\mathcal{U}=\left\{u_{1},u_{2},\ldots,u_{|\mathcal{U}|}\right\}$ denote the user set, $\mathcal{V}=\left\{v_{1},v_{2},\ldots,v_{|\mathcal{V}|}\right\}$ denote the item set, and $\mathcal{S}_{u}=\left\{v_{1},v_{2},\ldots,v_{n_{u}}\right\}$ denote the chronologically ordered interaction sequence for user $u\in\mathcal{U}$, where $n_{u}$ is the the length of the sequence. Given the interaction history $\mathcal{S}_{u}$, the task is to predict the next interacted item, denoted as $v_{n_{u+1}}$ for user $u$.
2.2. State Space Models
The State Space Model (SSM) is a sequence modeling framework based on linear ordinary differential equations. It maps an input sequence $x(t)\in\mathbb{R}^{D}$ to the output sequence $y(t)\in\mathbb{R}^{N}$ through the latent state $h(t)\in\mathbb{R}^{N}$:
(1) $h^{\prime}(t)$ $=\bm{A}h(t)+\bm{B}x(t),$
$y(t)$ $=\bm{C}h(t),$
where $\bm{A}\in\mathbb{R}^{N\times N}$ and $\bm{B},\bm{C}\in\mathbb{R}^{N\times D}$ are learnable matrices. To model discrete sequences instead of continuous functions, it must be discretized by a step size $\Delta$. Alternatively, the discretized SSM can be expressed as follows:
(2) $h_{t}$ $=\overline{\bm{A}}h_{t-1}+\overline{\bm{B}}x_{t},$
$y_{t}$ $=\bm{C}h_{t},$
where $\overline{\bm{A}}=\exp(\Delta\bm{A})$ and $\overline{\bm{B}}=(\Delta\bm{A})^{-1}(\exp(\Delta\bm{A})-\bm{I})\cdot\Delta\bm{B}$. After converting the parameters from a continuous form $(\Delta,\bm{A},\bm{B},\bm{C})$ to a discrete form $(\overline{\bm{A}},\overline{\bm{B}},\bm{C})$, the model can be computed in a linear recurrence way (gu2021combining), which enhances computational efficiency. Deriving from the vanilla SSM, the structured state space model (S4) (gu2021efficiently) imposes structure on the state matrix $\bm{A}$ with HiPPO (gu2020hippo) initialization to further improve the long-range dependency modeling.
Building upon S4, Mamba (gu2023mamba) introduces a data-dependent selection mechanism and leverages a hardware-aware parallel algorithm in recurrent mode. This combined approach empowers Mamba to capture contextual information effectively, particularly in long sequences, while maintaining computational efficiency. As a linear-time sequence model, Mamba achieves Transformer-quality performance with better efficiency, especially on long sequences.
3. Mamba4Rec
In this section, we introduce our proposed framework, Mamba4Rec. We begin with a high-level overview of the framework, followed by an exploration of its technical components. We describe how Mamba4Rec constructs a sequential recommendation model through an embedding layer, selective state space models, and a prediction layer. Additionally, we discuss the key components widely used in sequential recommendation, such as positional embeddings (vaswani2017attention), feed-forward networks, dropout (srivastava2014dropout), and layer normalization (ba2016layer).
3.1. Framework Overview
As illustrated in Figure 1, the proposed Mamba4Rec is a sequential recommendation model that leverages selective state space models through Mamba blocks. The core element of Mamba4Rec is the Mamba layer, which combines a Mamba block with a position-wise feed-forward network. Mamba4Rec offers flexibility in its architecture: while it can be stacked with multiple Mamba layers, a single Mamba layer often proves sufficient. Each layer comprises a Mamba block followed by a feed-forward network, enabling the model to capture both item-specific information and sequential context effectively from the user’s interaction history.
Figure 1. The overview of Mamba4Rec.
3.2. Embedding Layer
Similar to existing models, our approach utilizes an embedding layer to map item IDs to a high-dimensional space. This embedding layer uses a learnable embedding matrix $\bm{E}\in\mathbb{R}^{|\mathcal{V}|\times D}$, where $D$ is the embedding dimension. By applying the embedding layer to the input item sequence $\mathcal{S}_{u}$, we obtain the item embeddings $H$. To enhance robustness and prevent overfitting, we incorporate both embedding dropout and layer normalization after retrieving the embeddings:
(3) $H$ $=\operatorname{LayerNorm}\left(\operatorname{Dropout}\left(H_{o}\right)\right)\in\mathbb{R}^{n_{u}\times D}.$
During the training and inference stage, we merge multiple samples into one mini-batch, resulting the input data $x\in\mathbb{R}^{B\times L\times D}$ for the subsequent Mamba block, where $B$ is the batch size and $L$ is padded sequence length.
Positional Embedding. We investigate the impact of positional embeddings on our model’s performance. This includes evaluating learnable positional embeddings as used in (kang2018self; sun2019bert4rec). These embeddings are typically injected into the input item embedding before layer normalization. However, unlike Transformer-based models that rely on positional embeddings to capture the sequence order, state space models inherently handle sequential information through their recurrent nature. While theoretically unnecessary, we analyze the effect of positional embeddings on our model’s behavior in our experiments.
3.3. Mamba Layer
Table 1. Recommendation performance. The best results are bold, and the second-best are underlined.
Method MovieLens-1M Amazon-Beauty Amazon-Video-Games
HR@10 NDCG@10 MRR@10 HR@10 NDCG@10 MRR@10 HR@10 NDCG@10 MRR@10
NARM 0.2735 0.1506 0.1132 0.0627 0.0347 0.0262 0.1032 0.0530 0.0379
GRU4Rec 0.2934 0.1642 0.1249 0.0606 0.0332 0.0249 0.1030 0.0536 0.0380
SASRec 0.2977 0.1687 0.1294 0.0847 0.0425 0.0296 0.1168 0.0571 0.0390
BERT4Rec 0.3098 0.1764 0.1357 0.0760 0.0393 0.0282 0.1053 0.0538 0.0381
Mamba4Rec 0.3121 0.1822 0.1425 0.0812 0.0451 0.0362 0.1152 0.0603 0.0438
While the Mamba block (gu2023mamba) has potential as a sequence-to-sequence model, we leverage its strengths for sequential recommendation tasks. We introduce the Mamba layer, which incorporates both the Mamba block and a position-wise feed-forward network. Instead of simply stacking Mamba blocks, we use the Mamba layer as the fundamental building block for our architecture.
Mamba Block. As detailed in Algorithm 1, the Mamba block operates on an input $x_{t}\in\mathbb{R}^{B\times L\times D}$ with dimensions of batch size $B$, sequence length $L$, and hidden dimension $D$ at each time step $t$. It begins by performing linear projections on the input $x_{t}$ with an expanded hidden dimension of $D$ to obtain $x$ and $z$. These projections are then processed through a 1D convolution and a SiLU (elfwing2018sigmoid) activation. The core of the block involves a selective state space model with parameters discretized based on the input. This discretized SSM, along with $x^{\prime}$, generates the state representation $y$. Finally, $y$ is combined with a residual connection from $z$ after applying SiLU, and a final linear projection delivers the final output $y_{t}$ at time step $t$. Overall, the Mamba block leverages input-dependent adaptations and selective SSM to process sequential information effectively. The parameters of the Mamba block consist of an SSM state expansion factor $N$, a kernel size $K$ for the convolution, and a block expansion factor $E$ for input and output linear projections. Selecting the expansion factors $N$ and $E$ involves a tradeoff between (capturing complex relationships) and training efficiency (higher $N,E$ increase computational cost).
The Mamba block addresses the prevalent effectiveness-efficiency dilemma existing in RNNs and Transformers. The recurrent SSM in the Mamba block operates in linear time with respect to the sequence length, making it as efficient as RNNs for inference, achieved by discretized parameters. Additionally, the selection mechanism makes the parameters of SSM data-dependent. This enables Mamba to selectively remember or forget information, crucial for capturing complex dependencies within sequential data. With the structured state matrix addressing long-range dependencies with HiPPO, it outperforms self-attention blocks in modeling long sequences.
Algorithm 1 Mamba Block with Selective SSM
1:Input: $x_{t}:(B,L,D)$
2:Output: $y_{t}:(B,L,D)$
3:$x,z:(B,L,ED)\leftarrow\operatorname{Linear}(x_{t})$triangleright Linear projections of input
[…]
5:$\bm{A}:(D,N)\leftarrow\text{Parameter}$triangleright Structured state matrix
[…]
7:$\Delta:(B,L,D)\leftarrow\operatorname{Softplus}(\text{Parameter}+\operatorname{Broadcast}(\operatorname{Linear}(x^{\prime})))$
8:$\overline{\bm{A}},\overline{\bm{B}}:(B,L,D,N)\leftarrow\text{discretize}(\Delta,\bm{A},\bm{B})$triangleright Input-dependent parameters and discretization
9:$y:(B,L,ED)\leftarrow\text{SelectiveSSM}(\overline{\bm{A}},\overline{\bm{B}},\bm{C})(x^{\prime})$
10:$y^{\prime}:(B,L,ED)\leftarrow y\otimes\operatorname{SiLU}(z)$
11:$y_{t}:(B,L,D)\leftarrow\operatorname{Linear}(y^{\prime})$triangleright Final linear projection
12:return $y_{t}$
Position-wise Feed-Forward Network. We leverage a position-wise feed-forward network (PFFN) to improve the modeling of user actions in the hidden dimension:
(4) $\operatorname{PFFN}(H)=\operatorname{GELU}\left(H\bm{W}^{(1)}+\bm{b}^{(1)}\right)\bm{W}^{(2)}+\bm{b}^{(2)},$
where $\bm{W}^{(1)}\in\mathbb{R}^{D\times 4D},\bm{W}^{(2)}\in\mathbb{R}^{4D\times D},\bm{b}^{(1)}\in\mathbb{R}^{4D}$ and $\bm{b}^{(2)}\in\mathbb{R}^{D}$ are parameters of two dense layers and we use GELU (hendrycks2016gaussian) activation. The PFFN aims to capture complex patterns and interactions within the sequential data by applying two non-linear transformations using dense layers with the activation function. Additionally, to enhance model robustness and prevent overfitting, we utilize dropout and layer normalization, similar to Equation 3, after each Mamba block and feed-forward network within the Mamba layers. This helps to regularize the model and accelerate training convergence.
Stacking Mamba Layers We also explore using stacked Mamba layers in sequential recommendation. While deeper networks with multiple layers are not inherently beneficial for performance, residual connections (he2016deep) play a crucial role in facilitating the propagation of low-level features to higher layers throughout the multi-layer architecture. Recognizing the potential benefits of stacked Mamba layers, especially on long sequences, we explore two configurations for Mamba layers that balance effectiveness and efficiency:
  • Single Layer: When using only one Mamba layer, we apply layer normalization without residual connections. A single Mamba layer demonstrates comparable effectiveness to stacking Mamba layers while achieving higher efficiency in our tasks.
  • Stacked Layers: For stacked Mamba layers, residual connections become essential and enable the propagation of the last visited item’s embedding to the final layer.
To explore the tradeoff between effectiveness and efficiency, we further investigate the potential of stacked layers in experiments.
3.4. Prediction Layer
In the prediction layer, we adopt the approach of SASRec, using the last item embedding to generate the final output prediction scores:
(5) $\hat{y}=\operatorname{Softmax}\left(\operatorname{Linear}\left(h\right)\right)\in\mathbb{R}^{|\mathcal{V}|},$
where $h\in\mathbb{R}^{D}$ is the last item embedding from the Mamba layer. $\hat{y}\in\mathbb{R}^{|\mathcal{V}|}$ represents the probability distribution over the next item in the item set $\mathcal{V}$.
4. Experiments
4.1. Experimental Setup
Datasets. We conduct experiments and evaluate our models using three real-world datasets:
  • MovieLens-1M (harper2015movielens): A benchmark dataset contains about 1 million movie ratings from users.
  • Amazon-Beauty and Amazon-Video-Games (mcauley2015image): Two datasets contain product reviews and ratings for the “Beauty” and “Video Games” categories on Amazon.
For each user, we construct an interaction sequence by sorting their interaction records based on timestamps. We filter users and items with less than 5 recorded interactions, following the setting in previous works (kang2018self). The detailed statistics for each dataset after preprocessing are summarized in Table 2.
Table 2. Statistics of the experimented datasets.
Dataset # Users # Items # Interactions Avg. Length
ML-1M $6,040$ $3,416$ $999,611$ $165.4$
Beauty $22,363$ $12,101$ $198,502$ $8.9$
Video-Games $14,494$ $6,950$ $132,209$ $9.1$
Baselines. We compare Mamba4Rec against several baseline methods, which include RNN-based models like GRU4Rec (hidasi2015session) and NARM (li2017neural), as well as Transformer-based models such as SASRec (kang2018self) and BERT4Rec (sun2019bert4rec).
Evaluation Metrics. We adopt Hit Ratio (HR), Normalized Discounted Cumulative Gain (NDCG), and Mean Reciprocal Rank (MRR) with truncation at 10 as the evaluation metrics, i.e., HR@10, NDCG@10, and MRR@10.
Implementation Details. In the default architecture of our model, a single Mamba layer is employed without position embeddings. The Adam optimizer (kingma2014adam) is used with a learning rate 0.001. The training batch size is 2048, and the evaluation batch size is 4096. All models use an embedding dimension of 64. To address the sparsity of the Amazon datasets, a dropout rate of 0.4 is used, compared to 0.2 for MovieLens-1M. The maximum sequence length is set proportionally to the mean number of actions per user: 200 for MovieLens-1M and 50 for Amazon datasets. We follow RecBole (zhao2021recbole) for further implementation details. For the parameters of the Mamba block, the SSM state expansion factor is 32, the kernel size for 1D convolution is 4, and the block expansion factor for linear projections is 2.
4.2. Overall Performance
Table 1 demonstrates the overall recommendation performance. Among sequential recommendation methods, Transformer-based models SASRec and BERT4Rec generally outperform RNN-based baselines. Notably, the state-space-based Mamba4Rec achieves superior performance compared to both RNN and Transformer-based baselines across sparse and dense datasets, as well as varying maximum sequence lengths. Compared to SASRec, Mamba4Rec achieves significantly better performance on the MovieLens-1M dataset, with longer sequence lengths. This highlights the potential of state space models as a compelling alternative to Transformers and RNNs architectures for sequential recommendation.
4.3. Model Efficiency
Table 3 details the efficiency performance of Transformer-based models and Mamba4Rec, all measured on a single Nvidia A5000 GPU. Considering the tradeoff between efficiency and effectiveness, we use the model with a single Mamba layer instead of models with only a Mamba block or stacked Mamba layers on the long-sequence MovieLens-1M dataset. Mamba4Rec exhibits superior performance in terms of GPU memory usage, training time per epoch, and inference time per batch. Additionally, it demonstrates faster convergence, reaching its best validation performance around epoch 50, compared to SASRec (epoch 70) and BERT4Rec (epoch 180). These results suggest that Mamba4Rec achieves superior efficiency and reduces GPU memory cost, particularly on long-sequence datasets when compared to Transformer-based models.
Table 3. Efficiency performance on MovieLens-1M
Method GPU memory Training time Inference time
SASRec 14.98GB 131.24s 0.27s
BERT4Rec 15.48GB 207.60s 0.90s
Mamba4Rec 4.82GB 75.52s 0.14s
4.4. Ablation Study
To understand the contributions of each component in our architecture, we conduct an ablation study. Table 4 presents the results. We introduce each variant and analyze its respective impact:
  • Positional embedding: The results demonstrate that adding learnable positional embeddings does not improve the performance, due to the recurrent nature of the Mamba block.
  • Layer normalization: Following each Mamba block and feed-forward network, we employ dropout and layer normalization. Our results demonstrate their effectiveness in mitigating overfitting and enhancing performance.
  • Feed-forward network: The results indicate that the PFFN layer is particularly beneficial for the longer sequence dataset, MovieLens-1M. This suggests the PFFN’s ability to model nonlinearity leads to improvement in performance.
  • Mamba block only: Replacing the Mamba layer with a Mamba block (and removing PFFN, dropout, and layer normalization) results in faster computation but sacrifices overall performance.
  • Number of Mamba layers: Stacking two Mamba layers with residual connections yields performance improvements, particularly on the larger, long-sequence MovieLens-1M dataset. This shows the model’s ability to handle complex user behavior at the cost of reduced computational efficiency.
Table 4. Ablation analysis on ML-1M and Beauty.
Architecture Movielens-1M Amazon-Beauty
NDCG@10 MRR@10 NDCG@10 MRR@10
Default 0.1822 0.1425 0.0451 0.0362
Block Only 0.1786 0.1404 0.0405 0.0331
2 Layers 0.1845 0.1437 0.0440 0.0314
w/ PE 0.1738 0.1342 0.0408 0.0319
w/o PFFN 0.1740 0.1357 0.0437 0.0340
w/o LayerNorm 0.1820 0.1396 0.0419 0.0335
5. Conclusion and Future Work
In this work, we tackle the limitations of Transformer-based sequential recommendation models. We introduce Mamba4Rec built upon the selective structured state space model. Our experiments demonstrate that Mamba4Rec achieves strong performance on datasets with varying sparsity and sequence lengths. Additionally, it has significant improvements in both computational efficiency and memory cost, making it effective and efficient for sequential recommendation tasks. In the future, we aim to design state space models tailored for recommendation systems and promote the SSM-based models in this domain.